How much should Mrs. Dolores invest today in a time deposit with 5.5% interest if she expects to have P175, 000 for his son’s education at the end of 5 years?


1
Expert's answer
2021-11-16T14:21:05-0500

Let Mrs. Dolores investment today be represented by p.

A=p(1+r)t

p=A/(1+r)t

r=5.5%=0.055

A=175,000

t=5

p=175,000/1.0555

p= P 133,898.5119